ThatCreepyBoy:

Given the function f(x) = log2(x + 6), find the value of f−1(3). f−1(3) = 2 f−1(3) = 3 f−1(3) = 9 f−1(3) = 18

ThatCreepyBoy:

The first one :O

Vocaloid:

I don't have paper with me so I need to work this out here

ThatCreepyBoy:

Okie Dokie

Vocaloid:

y = log2(x+6) x = log2(y+6) 2^x = 2log2(y+6) 2^x = y + 6 y = 2^x - 6

Vocaloid:

yeah I get A as well

ThatCreepyBoy:

Okie thank you! I will be right back